OpenStudy (anonymous):

graph the equation y-3=-3/2(x+4) Please help

OpenStudy (zale101):

\(y-y_1=m(x-x_1)\) m is the slope <---you have this given (x1,y1) is the point. <--you have this given

OpenStudy (zale101):

Do you know how to plot a point in a Cartesian graph?

OpenStudy (anonymous):

no not really

OpenStudy (zale101):

Do you know what the term "slope" means?

OpenStudy (anonymous):

yes

OpenStudy (zale101):



OpenStudy (zale101):

Does this graph look familiar to you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need to graph the equation but i do not know how to

OpenStudy (zale101):

Do you know how to plot a point on this graph?

OpenStudy (zale101):

Yes, i'm walking you through on how tog graph on this plane.

OpenStudy (zale101):

Xy-plane

OpenStudy (zale101):

So, in your equation y-3=-3/2(x+4) What is \((x_1,y_1)\)?

OpenStudy (anonymous):

yes i know how to plot a point

OpenStudy (zale101):

Okay, good. Let us determine a point from your given equation y-3=-3/2(x+4).

OpenStudy (zale101):

@Hafsahilyas1 By looking at this, \(y-y_1=m(x-x_1)\) and looking at your equation \(y-3=-\frac{3}{2}(x-(-4))\) do you see any similarities?

OpenStudy (anonymous):

y1 is 3 and -3/2 is m and -4 ix x1

OpenStudy (zale101):

Perfect. Your point \((x_1,y_1)\) is \((3,-4)\). That's a point on the graph. Can you point \((3,-4)\) on this graph?